OpenStudy (anonymous):

Help please? Simplify Completely ((40v^2)/(35v^4)) + ((20v^3)/(5v))

OpenStudy (blurbendy):

I got 4v^2 + 8 / 7v^2

OpenStudy (mertsj):

\[\frac{40v^2}{35v^4}+\frac{20v^3}{5v}=\frac{8}{7v^2}+4v^2\]

OpenStudy (anonymous):

My options are: 2/7^4 7/32 7v^2/2 32/7

OpenStudy (anonymous):

?

OpenStudy (blurbendy):

Okay, then we can check our answer by substituting each of these answers in for v, and see which one gives us an equality.

OpenStudy (blurbendy):

Can you do that?

OpenStudy (mertsj):

You probably copied the problem wrong. Check it out.